The level of seismic activity during the last 24 hours has been low. There have been eighteen long-period earthquakes and a few large rockfall signals. Three volcano-tectonic earthquakes were located, at a depth of about 4 km beneath the crater area. The moderate amplitude broadband tremor on the Gages station ended at about 2100 last night.
EDM measurements were carried out this morning on the eastern triangle between White's, Long Ground and Castle Peak. No significant changes were measured.
During a helicopter observation flight this morning, good views of the crater area were obtained. The dome continues to steam vigorously, and some rockfalls were observed, which indicates continued dome growth. Yesterday afternoon, the dome was visible from the Observatory for the first time.